OpenStudy (anonymous):

Solve for x, y, z: x+y+z=12 6x-2y-z=16 3x+4y+2z=28

OpenStudy (anonymous):

HELP!!!!!

OpenStudy (anonymous):

what are you stuck on?

OpenStudy (anonymous):

solve the first equation for x and substitute into the last two equations... then solve those two equations for y and z

OpenStudy (anonymous):

I don't get any of it.

OpenStudy (anonymous):

"solve the first equation for x" move all the terms without 'x' in them to one side of the equal sign

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

I know THAT. x=-y-z+12

OpenStudy (anonymous):

great.

OpenStudy (anonymous):

6(-y-z+12)-2y-z=16

OpenStudy (anonymous):

"substitute into the last two equations." 6x-2y-z=16 - > 6(-y-z+12) -2y-z=16 3x+4y+2z=28 -> 3(-y-z+12) 4y+2z=28

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

do some basic algebra to simplify... what are you left with?

OpenStudy (anonymous):

two equations with two unknowns...

OpenStudy (anonymous):

solve them the same way we just 'solved' for x...

OpenStudy (anonymous):

solve one equation for z or y.... substitute into the other equation.

OpenStudy (anonymous):

what's16-72?

OpenStudy (anonymous):

good one.

OpenStudy (anonymous):

never mind.

OpenStudy (arhin):

x=4 y=0 z=8

OpenStudy (anonymous):

x=4,y=0 & z=8